OpenStudy (anonymous):

Fan and Medal!! The area of a rectangular piece of land is 240 square meters. If the length of the land was 5 meters less and the width was 3 meters more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. Part B: What is the width of the land in meters? Show the steps of your work.

OpenStudy (danjs):

A square would be X by X This rectangle is said to be (X-5) by (X+3)

OpenStudy (danjs):

(X-5)(X+3) = 240

OpenStudy (anonymous):

(x+3)(x-5)=240?

OpenStudy (anonymous):

Okay I got that

OpenStudy (danjs):

yes

OpenStudy (anonymous):

What do I do now?

OpenStudy (anonymous):

do you know how to find the area of a rectangle

OpenStudy (danjs):

Solve for X, you will get 2 answers, the positive one is correct

OpenStudy (anonymous):

do I FOIL (x+3)(x-5)?

OpenStudy (danjs):

yes

OpenStudy (anonymous):

I ended up getting x^2-2x-15

OpenStudy (danjs):

move the 240 over now to get x^2 - 2x - 255 = 0

OpenStudy (danjs):

factor that, or use the quadratic formula , since it is a book prob, it will probably factor nice

OpenStudy (anonymous):

Do I make that into two binomials?

OpenStudy (danjs):

yep (x-17)(x+15)=0

OpenStudy (anonymous):

Wait how did you get the 17 and 15?

OpenStudy (danjs):

255 = 17 * 15 and 15 - 17 = -2 (middle number)

OpenStudy (danjs):

If you multiply that out, you get back to the original ax^2 + bx + c = 0

OpenStudy (anonymous):

Ohhh okay I understand that

OpenStudy (danjs):

or you could always just apply the quadratic formula, and you would get the X values

OpenStudy (danjs):

so now you have (x-17)(x+15)=0 which means either term can be zero, so (x-17)=0 or (x+15)=0 X=17 or X=-15

OpenStudy (danjs):

The problem is a physical measure, so the positive value is the one that makes sense... X=17

OpenStudy (anonymous):

What do I write for part A and for Part B?

OpenStudy (danjs):

First Part... Area of Land = Length * Width = 240 240 = (X-5)(X+3)

OpenStudy (anonymous):

And do I work it out until I get (x-17)(x+15) = 0?

OpenStudy (danjs):

right, for the second part you solve it

OpenStudy (anonymous):

Thank you sososo much! You really helped a lot.

OpenStudy (danjs):

yw

OpenStudy (anonymous):

Do you think you can check over some of my answers? If you have the time of course

OpenStudy (danjs):

sure

OpenStudy (anonymous):

Thanks so much! They're quite long, so do you want me to screenshot it and attach it?

OpenStudy (danjs):

Sure, i have a bit of time

OpenStudy (anonymous):

And that's it, since I don't want to have you check all of them, haha

OpenStudy (danjs):

The first one looks good, what is the prob with that one?

OpenStudy (anonymous):

I just wanted to see if I got it correct or not

OpenStudy (danjs):

That one looks good to me

OpenStudy (anonymous):

Thank you (:

OpenStudy (danjs):

second one looks good too

OpenStudy (anonymous):

yay ^-^

OpenStudy (danjs):

For the last one part C, it asks if they all have a common factor

OpenStudy (anonymous):

Yes, is it wrong?

OpenStudy (danjs):

I don't see a common factor btw all 3 of those

OpenStudy (anonymous):

So it would be no because no common factor is consistent between the three polynomials?

OpenStudy (danjs):

yeah i think so

OpenStudy (anonymous):

Do you think you could check 2 more?

OpenStudy (danjs):

k, real quick, then i have to go... you can fan me for help later if you want

OpenStudy (anonymous):

Alright! Once I finish with these questions, I'll be done with Algebra for the year, aha

OpenStudy (danjs):

haha

OpenStudy (anonymous):

And that's it!

OpenStudy (danjs):

For the first, for the domain, be sure to call the variable 't' and not x, Since y is a function of t, in S(t)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

anything else?

OpenStudy (danjs):

nope, looks good to me

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

Did you check the last one? or

OpenStudy (danjs):

Part b asks what the maximum profit is, you dont say what that is

OpenStudy (anonymous):

How do I find the maximum?

OpenStudy (danjs):

it would be at the vertex of the parabola, or axis of symmetry value

OpenStudy (anonymous):

How do I find that? Did my steps correspond to the question at all?

OpenStudy (danjs):

It looks like you have P(n) written in vertex form in part b right... so you can use that to extract the vertex coordinate (h,k)

OpenStudy (anonymous):

Huh?

OpenStudy (danjs):

vertex point - (h,k) Vertex Equation - y = a(x-h)^2+k

OpenStudy (anonymous):

Do I need to rewrite anything?

OpenStudy (danjs):

so the vertex from your equation is at the point, (5, 2250)

OpenStudy (anonymous):

This is so confusing ._. I'm sorry that I'm so difficult to work with

OpenStudy (danjs):

sorry i said year... the most profit is at (5,2250) price 5 and prfit 2250

OpenStudy (anonymous):

How do I find that with the steps I wrote down?

OpenStudy (danjs):

which is also on the axis of symmetry

OpenStudy (danjs):

The last thing you wrote, is the vertex equation for the parabola. vertex point - (h,k) Vertex Equation - y = a(x-h)^2+k The maximum profit 'k' is at that vertex point, you can just read it from your equation.

OpenStudy (anonymous):

Ahhh okay! So now I write that the maximum profit is 2250 ?

OpenStudy (danjs):

yes, The vertex is at point (5,2250) When the tickets are sold at 5 dollars each, the profit is the most at 2250 dollars.

OpenStudy (anonymous):

Thank you sosososo much!! You're truly a lifesaver, I wish I could give you a thousand medals , haha

OpenStudy (danjs):